A simple generator has a square armature 6.0 cm on a side. The armature has 85 turns of 0.59-mm-diameter copper wire and rotates in a 0.800-T magnetic field. The generator is used to power a lightbulb rated at 12.0 V and 25.0 W. At what rate should the generator rotate to provide 12.0 V to the bulb? Consider the resistance of the wire on the armature.
Answer #1

give,

side = 6 cm

turns = 85

diameter of wire = 0.59 mm

magnetic field = 0.8 T

rating of bulb = 12 V and 25 W

resistivity of copper = 1.7 × 10-8 ohm m

resistance = resistivity * length / area

resistance = 1.7 × 10-8 * 0.06 * 4 * 85 / (pi * (0.59 * 10^-3 / 2)^2)

resistance of wire = 1.268 ohm

power = V * I

25 = 12 * I

I = 2.08 A

induced magnetic field = -N * B * A / t

2.08 * 1.268 = -85 * 0.8 * 0.06 * 0.06 * x / t

x/t = 10.77 rotation per second

rate should the generator rotate =  10.77 rotation per second